A stationary ball, with a mass of 0.5 kg, is struck by an identical ball moving at
6.0ms .
1
After the collision, the second ball moves 30° to the left of its original
direction. The stationary ball moves 60° to the right of the moving ball’s original
direction. What is the velocity of each ball after the collision?
1
Expert's answer
2020-03-23T10:55:02-0400

"\overrightarrow{v}" - velocity of the second ball before the collision

"|\overrightarrow{v}|=6\frac{m}{s}"

"\overrightarrow{v_1},\overrightarrow{v_2}" - velocities of the first and second balls after the collision accordingly



"m" - the mass of each ball

According to the law of impulse conservation,

"m\overrightarrow{v}=m\overrightarrow{v_1}+m\overrightarrow{v_2}\rArr \overrightarrow{v}=\overrightarrow{v_1}+\overrightarrow{v_2}"

So "\overrightarrow{v}" is the summ of vectors "\overrightarrow{v_1}" and "\overrightarrow{v_2} \rArr" according to the triangle law of adding vectors, "\overrightarrow{v}, \overrightarrow{v_1}, \overrightarrow{v_2}" form a triangle



The angle between "\overrightarrow{v_1}" and "\overrightarrow{v_2}" is equal to "30^o+60^o=90^o"


So this triangle is right "\rArr |\overrightarrow{v_1}|=|\overrightarrow{v}|cos60^o=6\frac{m}{s}*\frac{1}{2}=3\frac{m}{s}"


"|\overrightarrow{v_2}|=|\overrightarrow{v}|cos30^o=6\frac{m}{s}*\frac{\sqrt{3}}{2}=3\sqrt{3}\frac{m}{s}\approx5.2\frac{m}{s}"